IBM 1130 system restoration during 2011

Post this page to popular social media

Each TNMOC project has either a working group or project team assigned to do the work. Working groups are either managed in association with the CCS (Computer Conservation Society) or solely within the Museum.

The system at TNMOC is an IBM 1130 model 2B with 8KW 3.6us core store, internal IBM 2310 512KW removable disk drive, IBM 1132 printer and an IBM 1442 model 6 card reader/punch. We also have an IBM 29 Card Punch which is part of the TNMOC collection before the 1130 arrived. The 1130 system is on loan from The Museums of Liverpool and arrived at TNMOC on 22nd May 2009.

Below you can read the logs from the restoration project team of Stewart Baker (lead) and Peter Vaughan. This document is in reverse order of time with the most recent activities at the top. You can click on any of the images to see a full size version in a new window.

29/12/2011 Update from Peter Vaughan

Work continues with 1131 processor diagnostics.

A set of 7 one card diagnostic programs (ocdp) exists to test some of the basic instructions in the processor. As these cannot be loaded from the 1442, we are entering the instructions from assembly listings straight into core store using the console switches. So far ocdp01 -> 04 have been entered and run successfully which has tested a number of load/store/add/xor/wait and branch instructions. ocdp05 has been loaded but currently fails so further investigations are needed to discover why.

There appears to be a bit fault at core memory location 0x50 which is reproducible when reading the store in display mode, however it does not appear to fail the simple CE engineering core store write and read test so it is not clear if this is a fault or not.

28/12/2011 Update from Peter Vaughan

An initial audit of the punch cards that arrived with the 1130 has been done. We appear to have a master and working copy of a number of diagnostics card sets plus a number in a different draw labelled as old versions. Details of these still need to be documented.

We also have one copy only of a number of ripple card sets for testing alignment of the 1442.

Many of these diagnostics sets are too big to copy using the IBM 29 (which can only copy one card at a time) so some other method is needed to make working copies for us to use. Where possible we only want to work on our own copies of the cards.

Using a working copy of the 7 card set of one card diagnostics, a copy of each card was attempted using the IBM 29. After several attempts and a few crunched cards, copies were made of all but card 05, which keeps causing a card feed issue when a duplicate is made. During the duplication punching sometimes started at column 2 rather than 1. More work is needed to improve the reliability of the 29 and to discover why it does not like card 05.

Any attempt to load the cards into the 1131 from the 1442 results in the 1442 going into an error state so the only option is to enter the programs by hand.

22/12/2011 Update from Stewart Baker

Fault finding of IBM 1131 CPU started. System oscillator waveform examined and frequency read as 2.25MHz - this equates to a 3.6us store (the slower of the 2 types fitted).

Discovered that CE panel switches work in opposite sense to Test Lamp switch - there was no indication of correct position in any IBM manual or panel legend. It turns out that what we thought was the system running the last program in store when program start was pressed was actually the CPU running a store read/write test.

Now able to load and read from core memory. Simple routines used to load pattern into memory. FEMM 1.7.5 Core Store Console Isolation test run without any core parity errors.

17/12/2011 Update from Stewart Baker

IBM 29 Card Punch has an issue duplicating single cards. There is not enough skew adjustment at the read station due to the wear in a small bearing. Feeding cards to be duplicated through the normal card path works OK.

New Light Source bulb of the original type, bought on E-bay, fitted to the 1442, so no optical alignment should be necessary.

Hopper path solenoid back stop adjusted. Jam switch tape clearances reset. Mechanical adjustments made as per the IBM 1442 FEMM to items in the card path of the Stacker Cornering Station. A couple of these adjustments were a long way out. No repeat of card jams so far...

IBM 1131 CPU operation is not correct when trying to use the 1442 so further discussion needed on the best way to proceed with the fault finding.

13/12/2011 Update from Stewart Baker

Rough setting of Light Source voltage made (needs fine adjustment as per FEMM). Unit now going Ready with no Status errors.

Card path corrosion cleaned off with Brasso. Some pitting evident in aluminium, but should not effect operation of unit. Freed small seized bearing after punch station, and refitted loose metal guide piece. Unit now feeds cards, although it occasionally jams by selecting the wrong output hopper path (adjustment ?)

All local functions on IBM 1442 work, together with Program Load and Reset from the IBM 1130. We are close to trying to boot the IBM 1130 from punched cards...

08/12/2011 Update from Stewart Baker

Worked with Peter Vaughan to locate cause of IBM 1442 reported errors. Discovered Light Source Lamp blown. Lamp is a 8V 50W special made by Osram, however they are still available. One listed on E-bay.

Changed lamp for a spare, which although not the same shape has the same electrical characteristics.

Unit now giving READ STA and PUNCH STA errors. The READ STA error is due to Channel 12 of Read Head not working. This maybe due to either dirt or more likely insufficient light through that Fibre Optic channel.

Light Source Lamp adjustment needs checking, together with a clean on the Read Head assembly.

The card path has some corrosion of the aluminium, particularly at the Stacker Cornering Station. This should polish off.

06/12/2011 Update from Stewart Baker

New capacitor (4,700uF 63V) piggy backed across C4 to replace failed part in 1442.

IBM 1442 Card Reader/Punch connected to IBM 1130 with PSU cable and system powered on. No smoke or other form of distress observed. Internally generated +12V and 4.3V AC rails checked OK.

Data cable connected to IBM 1130 and system powered on. Again no smoke or other form of distress observed in either the 1442 or 1131.

Power On, Check and Trans (on status panel) lamps lit. The Check/Trans (Stacker Jam) error is not being caused by a card jam or defective micro switch. Need to investigate signal path as it may being generated in the 1131, however without circuit diagrams this will be difficult.

30/11/2011 Update from Stewart Baker

Continued working on the 1442 Card Reader/Punch - Internal clean carried out. Fuses and panel bulbs checked. Tightness of all screws on connector panels checked. Close visual examination of the PSU carried out. Contacts cleaned on switches and relays in the PSU. C2 and C3 (17,000uF 14V) capacitors reformed OK. C4 (4,000uF 15V) proved to be open circuit, replacement/substitute needed.

24/11/2011 Update from Stewart Baker

New roller micro switch fitted to IBM 29 card stack monitor to replace faulty part.

Interface to 1131 console printer disconnected, however it still performs un commanded CR/LF's.

16/11/2011 Update from Stewart Baker

IBM 29 card path adjustments made according to FEMM using Gram and Feeler gauges. All card punch functions including DUP working OK. Card punch now available for general use.

12/11/2011 Update from Stewart Baker

IBM 29 Card Punch: Cleaned card paths and removed metal burrs to enable consistent card feeding. Cards being fed, but no punch operation. Two micro switches in card path gummed up with grease and dirt - cleaned. Unit now punches cards correctly, but card feed from middle to left station unsatisfactory. Further mechanical attention needed.

02/11/2011 Update from Stewart Baker

New Drive Belt fitted to IBM 29. New spare and original left inside machine. Machine powered up. Basically working, but needs some maintenance.

Some keyboard keys stuck down. Result of 'must-touch-it-itis' ?

Investigation of 1131 console printer with Olly. Some progress, however the printer still sits there doing CR/LF's.

26/10/2011 Update from Stewart Baker

IBM 29 Card Punch cleaned internally. Mechanics lubricated with 3-in-one oil as per schedule in the FEMM, but re-greasing felt to be unnecessary.

Main motor Drive Belt removed as it has two splits in it, one almost through the belt to the canvas. Replacement belt sought.

25/10/2011 Update from Stewart Baker

IBM 29 Card Punch cleaned externally.

Blown bulbs in 1131 desk panel replaced with spares.

IBM 2315 disk cartridge examined and cleaned. The surface shows signs of contamination, but the biggest issue is a scratch on the lower surface in line with the index mark. This not bad enough to instigate a head crash, however the disk should only be used for testing, and not data storage.

Attempt to run up disk in 1131 was initially unsuccessful as the drive would not come up to speed (1500rpm). This prevented the heads from being loaded on to the disk surface. It was thought that lack of lubrication might be the cause, and so after consulting the FEMM for the drive, specific items were lubricated. These included the spindle drive motor which has a channel in its top casting for such a purpose.

After the motor was lubricated the heads loaded OK and the File Ready light came on. No head crashing or other undesirable effects occurred. The manual says that there is a 90 sec delay before the heads are loaded, but it was observed to be almost 2 mins.

The heads were examined and no signs of oxide deposit was found. Exercising the drive was left to a later date as it involves disconnecting data cables so that the internal CE switches become active.

19/10/2011 Update from Stewart Baker

Blown bulbs and ones with high resistance filaments changed in 1131 Status Panel. All bulbs now working.

Initial investigation of IBM 29 Card Punch - It appears to be in good condition. Unit cleaned internally, screw and paddle boards connections checked. Operating arm and roller missing from card stack micro switch. Should be easy to replace.

16/10/2011 Update from Stewart Baker

Web research shows that Selectric Typewriters/Printers that had been idle for a long time were immersed in a solvent bath before being re-lubricated. Although a maintenance manual for a Selectric Keyboardless Printer has been located, specialist assistance may be needed to restore the unit to a fully working condition.

12/10/2011 Update from Stewart Baker

Console Printer mechanics very stiff. Loosened/removed stale grease with copious amounts of WD40. Motor now drives the mechanics, with no more 'Hot smell', however it chatters like a Teletype all the time. Need to investigate whether it still does this when the interface to the 1131 is disconnected. There is a power switch underneath the printer to permit mains isolation without the need to power down the 1131.

23/09/2011 Update from Stewart Baker

Mulling over the PSU problem, looked at photo of MidPack regulators taken before their removal and spotted the link (or absence of !). Link fitted to +3V terminal block and system now powers up and stays up. '''!!! Whoopee !!!'''

Measured supply rails at test points on CPU chassis. All well within tolerance; -3V (-3.08V), +3V (+3.08), +6V (+6.1V), +12V (12.7V), +48V (+49.5V).

Display lamps tested. Only 3 bad. CPU Runs, Single Instruction and Single Steps (something left in core memory).

Mains input kept around 210V AC to reduce stress, and system left to soak for 2 hours.

'Hot smell' noticed. Eventually traced to stalled motor in Console Printer. The motor is required to run all the time the 1130 is powered on. Investigated Console Printer. Motor is working, but stalls trying to drive very stiff mechanics.

So, quite a lot of success today

21/09/2011 Update from Stewart Baker

1131 Power up task commenced - the first in over 25 years!

Unit moved into clear area, all covers removed and CPU & SDA racks opened. All PSU fuses removed and stored. All circuit breakers and power switches OFF.

Unit powered from Variac with an AVO in circuit to measure current. Negligible leakage current measured at 220V AC applied.

Circuit breaker CB1 closed and 20mA measured when relay R3 pulls in (24v AC T1). Fitted fuse F5 and closed Power and CE switch. Contactor K1 pulls in at approx 150V AC input.

All fans running, but system powers itself down after 5 seconds due to a missing MidPack PSU rail or rails. Fitted fuse F1 and closed breakers on MidPack PSUs to see if that was the problem. Found that MidPack Bulk PSU needs to be on to supply unregulated inputs to MidPack regulators.

Fitted fuse F2 and gradually powered system to 220V AC input. No new issues, but the system still powers down after 5 secs.

Measured continuities between MidPack Bulk PSU and MidPack regulator inputs all OK. Measured input voltages into MidPack regulators from Bulk PSU, all OK (approx 6.6V and 11.5V). Found +3V MidPack rail missing, all others OK. Swapping regulator PCB with -3V one did not cure the problem.

+3V MidPack PSU removed for bench diagnosis and rectification. Olly's help was much appreciated in power up activity.

After +3V PSU is fixed only the lamp supply and convenience outlets need checking as part of the 1131 power up actions.

14/09/2011 Update from Stewart Baker

Examination and cleaning of Bulk PSU - Distortion/discolouring of perspex shield over capacitors at the PSU rear investigated. This has been caused by heating from bleeder resistors. Shield heated and flattened. In an attempt to prevent further damage the shield has been positioned further from the resistors using washers as spacers.

Bulk PSU capacitors C1 to C9 checked for ESR, leakage and then reformed. All capacitors are in good physical and electrical condition. Bulk PSU refitted in 1131 chassis.

05/08/2011 Update from Stewart Baker

Continued 1132 line printer rust treatment on metal framework and front doors.

01/08/2011 Update from Stewart Baker

Continued 1132 line printer rust treatment on metal framework, side panels and rear doors. Side panels re hung. Repositioned IBM suite in display area.

05/07/2011 Update from Peter Vaughan

Bulk +12/+48v PSU removed after labelling all connectors. Visual inspection shows evidence of overheating on bleed resistors underneath plastic cover. PSU contains many high capacity capacitors which need to be reconditioned prior to initial testing of PSU.

We are close to wanting to power up the machine once all PSUs have been checked. A plan of action and risk assessment is being done to reduce the chances of damage to the machine should a PSU fail.

01/07/2011 Update from Stewart Baker**

Midpack PSU over voltage crowbar triggers measured and labelled. +3V (4.07V), -3V (4.17V), +6V (7.15V)

25/06/2011 Update from Stewart Baker

First sortie conducted on 1132 line printer. Tractor assembly de-rusted using electrolytic treatment. Initial stage of rust treatment to metal framework carried out. Equipment panels cleaned.

Disk cartridge removed from 2310 drive. Dust brushed and vacuum out of drive. R/W heads inspected using a mirror and light, no damage or oxide build up present. R/W heads cleaned with cotton buds and IPA, no oxide residue found on cotton buds. R/W head suspension springing examined, no defects found.

Surface of the 2315 disk cartridge examined. No evidence of damage found, however its surface is very dirty and needs cleaning with lint free cloths and IPA before any attempt is made to test the disk system.

12/06/2011 Update from Stewart Baker

Midpack Regulators -3v and +6v removed, cleaned and replaced in mainframe. As the +6v regulator card is easily removed, CAD documentation has been left till later. The over voltage crowbar PCB's need accurate checking of their trigger points.

All power supply fuses have been checked for continuity. Sticking fuse holders lubricated and cleaned.

The security of all accessible terminal block screws has been checked.

The internal cleaning of the 1131 is complete.

We are now subscribed to the mailing list. The '1130s Today' page on their web site contains details of our 1130 restoration, and links to TNMOC.

03/06/2011 Update from Stewart Baker

Electrolytic treatment of rust completed on 1131 Top covers (grey). Exposed / scratched areas on these panels treated with Carnauba wax. Exterior cleaning of 1131 now complete.

In order to better understand the Midpack Regulators functions, reverse engineering to CAD of the +3V unit has been carried out. This shows that in addition to the main DC supply into the regulator, there is another DC supply into the unit for the voltage regulator card. Like the main DC supply voltage it's value at this time is unknown.

The series regulator transistors are Motorola 108's which are germanium technology, and look like they were specifically made for IBM. They are also used in the LPT and are rare items. We appear to have no spares.

It is intended to finish the documenting of the +3V regulator before moving onto the -3V and +6V units. The -3V is identical to the +3V, however the +6V has some additional circuitry. These units will be cleaned and examined, however without detailed information being available (at present) apart from testing the Over voltage Crowbar protection PCB and looking for obvious faults, no powered testing outside the mainframe is practical.

Consideration has been given to removing and bench testing PS4 (+12v) and PS5 (+48v) located below the disk drive. It is my opinion that, because of their positioning the amount of upheaval and possible damage, that these supplies should be carefully tested in-situ. The 1131 system would appear to be well thought out in the provisioning of circuit breakers and fuses for the isolation and protection from fault conditions.

At present we have no schematic diagrams for any of the IBM 1130 suite, however there is a light at the end of the tunnel.. I have been successful in finally making contact with those running, and we are now in dialogue mode. In addition to publishing our details on the 1130 web site, they are looking to scan and provide us with the schematic documents we need.

13/05/2011 Update from Stewart Baker

Spare Selectric Console Printer retrieved from D-Block for evaluation and consideration as replacement for existing printer. Unfortunately this printer although in good condition having been refurbished in 1985 is not immediately suitable. a) Because it only supports pin feed paper b) Because it is much newer (model 55). The unit was photographed and returned to D-Block. Existing console printer exterior cleaned.

+3V PSU removed from 1131 and cleaned. Although no circuit diagrams are presently available the circuit seems quite simple. The unit is supplied with a smoothed DC from another PSU and only provides regulation and over voltage protection. Will be bench tested with external DC PSU and load.

Top covers (grey) removed from 1131 for Electrolytic treatment of rust.

27/04/2011 Update from Stewart Baker

All 1442 covers removed and cleaned. Most panels are in good condition; however there is bad scratch damage to upper rear panel. Dust brushed and vacuumed out.

Lamp and switch covers cleaned on 1131, 1132 and 1442.

1131 keyboard cover removed and the area cleaned. There is some surface rust present.

Filter under Power Supply Blower YP111 refitted with new filter medium.

Emergency Stop switch reset and tested.

20/04/2011 Update from Stewart Baker

Midpack Regulators MPR1,2 & 3 area cleaned and photographed. Tuned 115v/230v transformer cleaned and photographed. Front of Bulk Power Supplies (PSDC) cleaned and photographed.

Power Supply Blower (YP111) removed and cleaned. Fan in YP111 is 230v 50Hz. Fan run for 4 hours without problem. Filter under YP111 missing. Spare in Archive has a deteriorated filter medium - to be replaced.

Fuses in Power Sequence Box removed and checked for continuity.

Cream disk surround removed and cleaned. One of the grey plastic supports is missing, causing the switch panel to float.

15/04/2011 Update from Stewart Baker

External cleaning of 1131 completed. Most panels are OK, however the blue panel in front of the disk drive has rust under the paint. This is worst along the bottom edge of the panel. The punched card has been removed from the locking mechanism of the desk. This has now allowed the desk to fit without warping. The locking mechanism has been lubricated and can be operated using a screwdriver blade through the slot in a panel below the keyboard.

09/04/2011 Update from Stewart Baker

Meeting with Peter Vaughan to discuss all aspects of 1130 restoration.

On-site spares identified, together with those Peter is protecting at home.

Archived punch cards need to be re-labelled before the pencil markings fade completely.

Documation card reader found. Needs to be moved to the same area as the spare Selectric printer.

Need to find power distribution diagrams of the system before proceeding with PSU evaluation.

07/04/2011 Update from Stewart Baker

Start of Stewart Baker's involvement in the IBM 1130 restoration.

Equipment moved to interim positions for display/restoration.

First level external cleaning of the equipment carried out. The top cover of the Line Printer needed Meths to get out the ingrained carbon. The majority of panels should benefit from further cleaning using a variety of methods. The most obvious blemishes are the rusty scratches on the grey panels to the rear of the 1130 console. These will require attention.

All covers removed from the 1130 to view the internals. There is little sign of rust, and no obvious damage of any parts.

There is a potential hazard of getting ribbon cables trapped whenever the processor rack is swung in/out and great care must be taken to prevent any cable damage. My recommendation is that the CPU section should be accessed as little as possible.

The top cover was removed from the top of the disk drive to examine the condition of the heads, there having been a report that they were stuck to the removable drive platter. It was found (to much relief) that the heads were in the home position, and that no damage to the disk oxide surface was apparent. Further the spindle bearings, belts and motor show no signs of mechanical binding. The disk has been left in the mechanically locked position as a precaution against any potential damage.

The unit shows a considerable British content with labels bearing company names with 'Ltd'. In fact there is one from PYE. Most of these are attached to cable forms. Various paper labels that were in danger of becoming detached from the metalwork were temporarily secured at the top and bottom with tape.

To make future access to all areas of the machine easier, all hinges and locking pins were cleaned of corrosion and lightly lubricated with WD40.

Support us

The Museum has not received government or Lottery funding, so your help is needed.

Become a member »
Make a donation »
Become a volunteer »
Sponsor us »

Latest Tweets